OpenStudy (anonymous):

Solve sinx + cosx = 1 on 0

OpenStudy (loser66):

square both sides, what do you get?

OpenStudy (anonymous):

sin^2(X) +2sin(x)cos(x)+ cos^2(x) =1

OpenStudy (anonymous):

ohhh 1+sin(2x) =1

OpenStudy (loser66):

ohhh no

OpenStudy (loser66):

don't go that way, 2sinx cosx =0 then either sinx =0 or cos x =0. solve for them.

OpenStudy (anonymous):

oh got it , so its's 0 and pi/2?

OpenStudy (loser66):

pi/2, pi, 3pi/2,

OpenStudy (loser66):

on your problem \(0\leq x\leq 2\pi\) or just \(0<x<2\pi\)?

OpenStudy (loser66):

equal or not?

OpenStudy (anonymous):

0≤x≤2π

OpenStudy (loser66):

ok, so take 0 and 2pi also

OpenStudy (loser66):



OpenStudy (anonymous):

oh got it, thanks! :)